The sum of the differences between sample observations and the sample mean is equal to _______.

The range

Zero

The mean deviation

The standard deviationv
Answer #1

Zero because all deviations from sample mean add to 0

$$ \sum\left(\bar{x}-x_{i}\right)=n * \bar{x}-\sum x_{i}=\sum x_{i}-\sum x_{i}=0 $$